# CoxSE: Exploring the Potential of Self-Explaining Neural Networks with Cox Proportional Hazards Model for Survival Analysis

### v1

@@ Abstract
The Cox Proportional Hazards (CPH) model has long been the preferred survival model for its explainability. However, to increase its predictive power beyond its linear log-risk, it was extended to utilize deep neural networks sacrificing its explainability. In this work, we explore the potential of self-explaining neural networks (SENN) for survival analysis. we propose a new locally explainable Cox proportional hazards model, named CoxSE, by estimating a locally-linear log-hazard function using the SENN. We also propose a modification to the Neural additive (NAM) models hybrid with SENN, named CoxSENAM, which enables the control of the stability and consistency of the generated explanations.
Several experiments using synthetic and real datasets have been performed comparing with a NAM-based model, DeepSurv model explained with SHAP, and a linear CPH model. The results show that, unlike the NAM-based model, the SENN-based model can provide more stable and consistent explanations while maintaining the same expressiveness power of the black-box model. The results also show that, due to their structural design, NAM-based models demonstrated better robustness to non-informative features. Among these models, the hybrid model exhibited the best robustness.
@@ 1 Introduction
Due to their flexibility and the abundance of collected data, neural-network-based models have been shown to outperform classical machine learning methods in survival modeling [1]. However, such performance gain comes at the price of their explainability. Neural networks have long been considered black-box models because of their complicated inner workings. This makes it difficult to understand how they arrive at their predictions. This hinders the usability of such models especially in domains like survival analysis when used for safety-critical applications like some areas in health care or predictive maintenance [2, 3, 4, 5]. Nonetheless, efforts have been spent utilizing post-hoc explanation methods to compensate for the lack of transparency in black-box models [6].
@@
Recently, there has been a growing interest in a type of neural network that not only predicts the outcome of an instance but also provides inherent explanations of how it arrived at such an outcome. Examples of such an approach are the Neural Additive Models (NAM) [7] and the Self-Explainning Neural Networks (SENN) [8]. Such models can learn to fit the problem while providing intrinsic explanations for their decisions. The NAM method converts the problem into a sum of single-input functions of individual features. Each single-input function is modeled as a neural network, the output of which represents the respective feature’s contribution to the decision. On the other hand, SENN models the function as locally-linear function where the model learns the relevance of the features conditioned on the input.
@@
The NAM structure was first used for survival analysis in SurvNAM [9] as a surrogate Cox-based model to explain black-box survival models. Subsequently, it was recently proposed as a stand-alone model named CoxNAM [10]. However, due to its structural design, the NAM models cannot model feature interactions limiting their predictive power. Moreover, the NAM models provide explanations that highlight the contribution of each feature to the decision omitting the model’s sensitivity to change in those features.
@@
Therefore, in this work, we explore the potential of the SENN structure for survival modeling. We propose CoxSE, a Cox-based self-explaining model relying on the SENN structure. CoxSE learns features’ relevance which can be interpreted as the weights of the line tangent to the decision boundary in the vicinity of the point of interest. As all features contribute to the computation of each feature’s relevance, CoxSE can account for feature interactions. Additionally, CoxSE has two extra regularization terms to encourage the stability and consistency of the provided explanations.
@@
However, such flexibility comes with a price, as the NAM structure is shown to be more robust to noise. Therefore, we also propose a hybrid model CoxSENAM, adopting the NAM structure and the type of output of the SENN along with its loss function’s regularization terms. Similar to CoxNAM, such a model is more suitable for cases with no feature interactions, however, it benefits from the extra regularization of the learned features’ relevance, achieving better stability, consistency, and robustness.
@@ 2 Survival Analysis
Survival analysis is a branch of statistics concerned with studying the time until an event of interest occurs, for example, a patient’s death or a machine’s failure. The main challenge that gave rise to such a branch of science is that the collected data contains subjects that have not experienced the event yet at the end of the study. Such cases are called censored cases which are assumed to experience the event at a later time greater than their recorded time, in which case it is called right censored data. There are other types of censoring, however, right censoring is the most common.
Survival data usually comes as tuples of three variables ($\mathbf{x}_{i},t_{i},e_{i}$) that describe, for each subject ($i$), the features’ vector $\mathbf{x}_{i}$, the recorded time $t_{i}$, and the event indicator $e_{i}$ to indicate whether the recorded time is the time-to-event or the censoring time.
@@
The hazard function represents the instantaneous death rate which quantifies the risk of the event in the infinitesimal time $\delta t$ given that the subject has survived until the time $t$ and is defined as:
@@
Those two functions are related where:
@@
The cumulative hazard function (CHF) is another function of interest to some models and is related to the hazard and survival functions:
[…]
and:
@@
Early on, statistical methods have been developed to estimate such functions as the Kaplan-Meier estimator [11] to estimate the survival function or the Nelson-Aalen estimator [12, 13, 14] to estimate the CHF. However, the Cox Proportional Hazards Model (CPH) [15] is the first model to account for feature dependency. The CPH model assumes a baseline hazard function $h_{0}(t)$ common to all subjects and only depends on time. Whereas, the features $\mathbf{x}$ have exponential multiplicative effects on the hazard that is independent of time. As a result, there is a linear relationship between the features and the logarithm of the hazard function:
$h(t,\mathbf{x})=h_{0}(t)e^{\mathbf{w}^{\intercal}\mathbf{x}}$ (6)
Due to its linearity, the CPH model is considered to be explainable, where the weights $\mathbf{w}$ indicate the relative importance of the features in the decision.
Subsequently, machine learning started to be used for survival analysis, offering more accurate and flexible models. Also, more recently, with the abundance of collected data, deep learning models in particular emerged as the predominant approach. Numerous deep-learning models have been proposed to estimate different survival analysis functions employing various techniques and objective functions for this purpose [16, 17, 1, 18]. Interestingly, over $40\%$ of such models employ the Cox-based formulation [19]. Most notably DeepSurv [16] which extends the CPH linear log-risk function ($\mathbf{w}^{\intercal}\mathbf{x}$ in Equation 6) to a non-linear function ($f(x)$ in Equation 7) modeled as a neural network.
$h(t,\mathbf{x})=h_{0}(t)e^{f(\mathbf{x})}$ (7)
There are many evaluation metrics used to evaluate survival models. However, the concordance index (C-index) [20] is arguably the most frequently used metric. The C-index quantifies the rank correlation between the predicted scores of subjects and their observed or censored times. More specifically, it represents the probability that the predicted risk scores of two randomly selected subjects agree with the order of their observed event times as stated in Equation 8. Thus, a good model should assign higher risk scores for subjects with shorter survival times leading to a higher C-index value.
$CI=\mathbb{P}(\hat{r}_{i}>\hat{r}_{j}|t_{i}<t_{j},e_{i}=1)$ (8)
where $\hat{r}_{i}$ is the predicted risk score of the subject $i$ with the time-to-event value $t_{i}$ and event indicator $e_{i}$.
@@ 3 Related work
As machine learning started to be used for survival analysis, a need for suitable explainability methods emerged—-survival models usually output functions, which hinders the use of machine learning explanation methods. Many post-hoc explanation methods in machine learning like LIME [21] and SHAP [22] have been adapted to survival analysis models as in [23, 24, 25]. Moreover, other methods relied on counterfactual examples adapting them for survival models [26, 27].
However, a growing direction focuses on building inherently explainable, yet expressive, models. SurvNAM [9] proposed the use of Neural Additive Models (NAM) [7] and an extended Cox-Proportional Hazards model as a surrogate model to explain black-box survival models. More recently, CoxNAM [10] proposed using the NAM structure with the Cox formulation as a standalone inherently explainable survival model.
Originally, the NAM models the relation between an input $\mathbf{x}=(x_{0},x_{1},...,x_{p})$ and a target $y$ as an additive function based on the Generalized Additive Models (GAM) model [28]:
@@
Due to their design, NAM-based models cannot model interactions between features. This limits the expressive power of such models leading to underfitting in most cases. Moreover, there is no guarantee that the model will learn valid explanations.
Our work relies on another interesting direction for inherent explanations, the Self-Explaining Neural Networks (SENN) [8]. Such a model has two branches: a concept encoder that generates interpretable features $h(\mathbf{x})$ from the input and an input-dependent parameterizer that learns the features’ relevance scores (weights) $\mathbf{w}(\mathbf{x})$. Eventually, the function can be written as:
$f(x)=g(w_{1}(\mathbf{x})h_{1}(\mathbf{x}),..,w_{p}(\mathbf{x})h_{p}(\mathbf{x}))$ (10)
where $g(.)$ is the aggregation function and can be any affine function with positive weights or simply an addition. For tabular data, the concept vector is the original feature space, i.e. $h(\mathbf{x})=\mathbf{x}$, approximating the function with a locally-linear function. Additionally, SENNs regularize their generated explanations to ensure that similar subjects have similar explanations, a quality referred to as stability.
As the weights of each feature depend on all the features, SENN can model the interactions between the features maintaining similar performance as regular neural networks. Moreover, SENN with its regularization can control the quality of the explanations.
@@ 4 Method
Our method CoxSE extends the self-explaining neural network (SENN) [8] to survival analysis using the Cox Proportional Hazards model (CPH) [15], however, with non-linear relationship $f(\mathbf{x})$ between the predictors and the hazard as shown in Equation 11:
$h(t,\mathbf{x})=h_{0}(t)e^{f(\mathbf{x})}$ (11)
where $h_{0}(t)$ is an unspecified baseline hazard function and the log-risk function $f(\mathbf{x})$ is approximated as a locally linear function:
$f(\mathbf{x})=\sum_{x_{i}\in\mathbf{x}}w_{\theta_{i}}(\mathbf{x})x_{i}$ (12)
This allows the model to learn the relevance vector $\mathbf{w}_{\theta}(\mathbf{x})$ that specifies the linear weights at each point $\mathbf{x}$. Such weights describe a vector perpendicular to the decision boundary and reflect the local features’ importance at each specific location. An illustration of the CoxSE model is shown in Figure 2.
Figure 2: CoxSE Model architecture
Based on the SENN and the NAM models we also propose a hybrid mode named CoxSENAM. As illustrated in Figure 3, the model has a structure similar to the CoxNAM model and a type of output and loss function similar to the CoxSE model. In this model, features’ relevance are computed independently:
$f(\mathbf{x})=\sum_{x_{i}\in\mathbf{x}}w_{\theta_{i}}(x_{i})x_{i}$ (13)
where $w_{\theta_{i}}$ is the relevance function of the i’th feature in the features’ vector $\mathbf{x}$.
Figure 3: CoxSENAM Model architecture
The loss function consists of three terms. The first term is the Partial Likelihood function inherited from the CPH model, Equation 14. This part of the loss function maximizes the concordance between the predicted hazards.
$\mathcal{L}_{1}(\theta)=\prod_{i=1}^{N}\frac{e^{f_{\theta}(\mathbf{x}_{i})}}{\sum_{\mathbf{x}_{j}\in R_{i}}e^{f_{\theta}(\mathbf{x}_{j})}}$ (14)
@@
The second term, Equation 15, minimizes the distance between $\nabla_{x}f_{\theta}(\mathbf{x})$, the gradient of $f_{\theta}(\mathbf{x})$ with respect to $\mathbf{x}$, and the predicted local weights $w_{\theta}(\mathbf{x})$:
$\mathcal{L}_{2}(\theta)=\|\nabla_{x}f_{\theta}(\mathbf{x})-w_{\theta}(\mathbf{x})\|_{2}$ (15)
This term is adopted from the SENN model to encourage explanations’ stability. We also added a third term to the loss function that regularizes the learned weights and encourages the model to select a subset of the most important features while setting the weights of less important features to zero:
$\mathcal{L}_{3}(\theta)=\|w_{\theta}(\mathbf{x})\|_{1}$ (16)
The three terms comprise the loss function:

### v2

@@ Abstract
The Cox Proportional Hazards (CPH) model has long been the preferred survival model for its explainability. However, to increase its predictive power beyond its linear log-risk, it was extended to utilize deep neural networks, sacrificing its explainability. In this work, we explore the potential of self-explaining neural networks (SENN) for survival analysis. We propose a new locally explainable Cox proportional hazards model, named CoxSE, by estimating a locally-linear log-hazard function using the SENN. We also propose a modification to the Neural additive (NAM) model, hybrid with SENN, named CoxSENAM, which enables the control of the stability and consistency of the generated explanations.
Several experiments using synthetic and real datasets are presented, benchmarking CoxSE and CoxSENAM against a NAM-based model, a DeepSurv model explained with SHAP, and a linear CPH model. The results show that, unlike the NAM-based model, the SENN-based model can provide more stable and consistent explanations while maintaining the predictive power of the black-box model. The results also show that, due to their structural design, NAM-based models demonstrate better robustness to non-informative features. Among the models, the hybrid model exhibits the best robustness. Full implementation is available on GitHub. [^1] https://github.com/abdoush/CoxSE
@@ 1 Introduction
Due to their flexibility and the abundance of collected data, deep-neural-network-based models have been shown to outperform classical machine learning methods in survival modeling [ALABDALLAH2024]. However, such performance gain comes at the price of their explainability. Neural networks have long been considered black-box models because of their complicated inner workings. This makes it difficult to understand how they arrive at their predictions. Consequently, the black-box nature limits the usability of such models, especially in domains like survival analysis when used for safety-critical applications like health care or predictive maintenance [Reza:2021, Polymeri:2020, Alabdallah:2023Premature, Rahat:2023]. Nonetheless, efforts have been spent utilizing post-hoc explanation methods to compensate for the lack of transparency in black-box models [pashami:2023Xpm].
@@
Recently, there has been a growing interest in a type of neural network that not only predicts the outcome of an instance but also provides inherent explanations of how it arrived at such an outcome. Examples are the Neural Additive Models (NAM) [Agarwal2021] and the Self-Explaining Neural Networks (SENN) [Alvarez2018]. Such models can learn to fit the problem while providing intrinsic explanations for their decisions. The NAM method converts the problem into a sum of single-input functions of individual features, modeled as a neural network. The output of these functions represents the respective feature’s contribution to the decision. The SENN models the function as a locally-linear function where the model learns the relevance of the features conditioned on the input.
@@
The NAM structure was first used for survival analysis in SurvNAM [Utkin2022] as a surrogate Cox-based model to explain black-box survival models. Subsequently, it was recently proposed as a stand-alone model named CoxNAM [Xu2023]. However, owing to their structural design, NAM models are unable to capture interactions between features, thereby constraining their predictive capacity. Moreover, the NAM models provide explanations that highlight the contribution of each feature to the decision, omitting the model’s sensitivity to changes in those features. In contrast, SENN-based models are not constrained by limitations in modeling feature interactions and can attain predictive performance comparable to that of black-box models. Additionally, these models learn the relevance of individual features, reflecting their local sensitivities, from which the contributions of features to specific predictions can be derived.
@@
In this work, we explore the potential of the SENN structure for survival modeling. We propose CoxSE, a Cox-based self-explaining model relying on the SENN structure. CoxSE learns features’ relevance, which can be interpreted as the weights of the line tangent to the decision boundary in the vicinity of the point of interest. As all features contribute to the computation of each feature’s relevance, CoxSE can account for feature interactions. Additionally, CoxSE has two extra regularization terms to encourage the stability and consistency of the provided explanations.
@@
However, such flexibility comes with a price, as we demonstrate that the NAM structure is more robust to noise. Therefore, we also propose a hybrid model, CoxSENAM, adopting the NAM structure with the SENN type of output and loss function. Similar to CoxNAM, this model is more suitable for cases with no feature interactions, but benefits from the extra regularization of the learned features’ relevance, achieving better stability, consistency, and robustness to noise.
In summary, this paper provides three contributions: 1) proposing CoxSE, a new Cox-based survival model providing intrinsic explanations while maintaining the similar performance of its black-box counterpart; 2) proposing CoxSENAM, an enhancement to the NAM-based model adopting the SENN type of output and loss function which shows better explanation stability, consistency, and robustness to noise; 3) illustrating the pros and cons of the proposed and existing models.
@@ 2 Survival Analysis
Survival analysis is a branch of statistics concerned with studying the time until an event of interest occurs—for example, a patient’s death or a machine’s failure. The main challenge that gave rise to survival analysis is that survival data contains subjects who have not experienced the event yet at the end of the study. Such cases are called censored cases, which are assumed to experience the event at a later time greater than their recorded time, in which case it is called right-censored data. Other types of censoring exist, but right censoring is the most common.
Survival data usually comes as tuples of three variables ($\bm{x}_{i},t_{i},e_{i}$) that describe, for each subject ($i$), the feature vector ($\bm{x}_{i}$), the recorded time ($t_{i}$), and the event indicator ($e_{i}$) to indicate whether the recorded time is the time-to-event or the censoring time.
@@
The hazard function represents the instantaneous death rate, which quantifies the risk of the event in the infinitesimal time $\Delta t$ given that the subject has survived until the time $t$ and is defined as
@@
The two functions are related through the relation
@@
The cumulative hazard function (CHF) is another function of interest to some models and is related to the hazard and survival functions by
$H(t)=\int_{0}^{t}h(\tau)d\tau$ (4)
and
@@
Early on, statistical methods were developed to estimate such functions as the Kaplan-Meier estimator [Kaplan:1958] for the survival function, and the Nelson-Aalen estimator [Nelson:1969, Nelson:1972, aalan:1978] for the CHF. However, these methods do not account for feature dependencies. The Cox Proportional Hazards Model (CPH) [Cox1972] was the first statistical model to incorporate feature dependence in the analysis of time-to-event data. The CPH model assumes a baseline hazard function $h_{0}(t)$, which is common to all subjects and solely dependent on time, while the covariates $\bm{x}$ exert time-independent, multiplicative effects on the hazard through an exponential function. Consequently, the model imposes a linear relationship between the covariates and the logarithm of the hazard function:
$h(t,\bm{x})=h_{0}(t)e^{\bm{w}^{\intercal}\bm{x}}$ (6)
Due to its linearity, the CPH model is regarded as inherently explainable, with the weights $\bm{w}$ representing the relative importance of the features in influencing the hazard.
Over time, machine learning techniques have been increasingly applied to survival analysis, enabling the development of more accurate and flexible models. More recently, the widespread availability of large-scale data has led to the emergence of deep learning models as the dominant approach in this domain. Numerous deep-learning models have been proposed to estimate different survival analysis functions, employing various techniques and objective functions for this purpose [Katzman:2018, Lee_2018, ALABDALLAH2024, Altarabichi:2024Loss]. Interestingly, more than $40\%$ of such models employ the Cox-based formulation [Wiegrebe2024]. Most notably, DeepSurv [Katzman:2018] which extends the CPH linear log-risk function ($\bm{w}^{\intercal}\bm{x}$ in Eq. 6) to a non-linear function ($f(\bm{x})$ in Eq. 7) modeled as a neural network.
$h(t,\bm{x})=h_{0}(t)e^{f(\bm{x})}$ (7)
With the growing variety and complexity of survival models, robust and consistent evaluation becomes essential. Among the many metrics available, the concordance index (C-index) [Harrell:1982] is arguably the most widely used. The C-index (CI) quantifies the rank correlation between the predicted scores of subjects and their observed or censored times. More specifically, it represents the probability that the predicted risk scores of two randomly selected subjects agree with the order of their observed event times:
$\mbox{\emph{CI}}=\mathbb{P}(\hat{r}_{i}>\hat{r}_{j}|t_{i}<t_{j},e_{i}=1)$ (8)
where $\hat{r}_{i}$ is the predicted risk score of the subject $i$ with the time-to-event value $t_{i}$ and event indicator $e_{i}$. A good model should assign higher risk scores for subjects with shorter survival times, leading to a higher C-index value.
@@ 3 Related work
3.1 Post-Hoc Explainability Methods
As machine learning started to be used for survival analysis, a need for suitable explanation methods emerged. Many post-hoc explanation methods in machine learning like LIME [Ribeiro_2016] and SHAP [Lundberg_2017] have been adapted to survival analysis models as in [KOVALEV2020, Alabdallah2022, KRZYZINSKI2023, Langbein_2025, Baniecki_2025]. Moreover, other methods relied on counterfactual examples, adapting them for survival models [Kovalev:2021Counterfactuals, Alabdallah:2024Counterfactuals].
3.2 Inherently Explainable Models
A growing research direction focuses on building inherently explainable, yet expressive, models. SurvNAM [Utkin2022] proposed the use of Neural Additive Models (NAM) [Agarwal2021] and an extended Cox-Proportional Hazards model as a surrogate model to explain black-box survival models. More recently, CoxNAM [Xu2023] proposed using the NAM structure with the Cox formulation as a standalone inherently explainable survival model.
Originally, the NAM models the relation between an input $\bm{x}=(x_{0},x_{1},...,x_{p})$ and a target $y$ as an additive function based on the Generalized Additive Models (GAM) model [Hastie1986]:
@@ 3.2 Inherently Explainable Models
Due to their design, NAM-based models cannot model interactions between features. This limits the expressive power of such models, leading to underfitting in most cases. Moreover, there is no guarantee that the model will learn valid explanations.
Our work builds on another interesting direction for inherent explanations: the Self-Explaining Neural Networks (SENN) [Alvarez2018]. Such a model has two branches: a concept encoder that generates interpretable features $\bm{h}(\bm{x})$ from the input and an input-dependent parameterizer that learns the features’ relevance scores (weights) $\bm{w}(\bm{x})$, as illustrated in Figure 2.
Figure 2: SENN Model architecture
Eventually, the function can be written as:
$f(x)=g(w_{1}(\bm{x})h_{1}(\bm{x}),..,w_{p}(\bm{x})h_{p}(\bm{x}))$ (10)
where $p$ is the number of features and $g(.)$ is the aggregation function, which can be any affine function with positive weights or simply an addition, approximating the function with a locally-linear function as:
$f(x)=\sum_{x_{i}\in\bm{x}}w_{i}(\bm{x})h_{i}(\bm{x})$ (11)
Additionally, SENNs regularize their generated explanations to ensure that similar subjects have similar explanations, a quality referred to as stability.
For tabular data, which is the case in our work, the concept vector is the original feature space, i.e. $\bm{h}(\bm{x})=\bm{x}$. This eliminates the need for the autoencoder branch of the SENN structure, along with its reconstruction part of the loss function.
As the weights of each feature depend on all the features, SENN can model the interactions between the features, maintaining similar performance as regular neural networks. Moreover, SENN, with its regularization, can control the quality of the explanations.
@@ 4 Method
In this work, we investigate the use of Self-Explaining Neural Networks (SENN) and Neural Additive Models (NAM) to improve interpretability in survival analysis. We adapt the SENN framework to the Cox Proportional Hazards (CPH) model and integrate key elements from both SENN and NAM to develop two novel models: CoxSE and CoxSENAM. Our aim is to provide locally interpretable risk predictions while retaining the flexibility of non-linear models.
4.1 Self-Explaining Cox Model (CoxSE)
The CoxSE method extends SENN to survival analysis using the CPH model, but with a non-linear relationship $f(\bm{x})$ between the predictors and the hazard:
$h(t,\bm{x})=h_{0}(t)e^{f(\bm{x})}$ (12)
where $h_{0}(t)$ is an unspecified baseline hazard function and the log-risk function $f(\bm{x})$ is approximated as a locally linear function
$f(\bm{x})=\sum_{x_{i}\in\bm{x}}w_{\theta_{i}}(\bm{x})x_{i}$ (13)
This allows the model to learn the relevance vector $\bm{w}_{\theta}(\bm{x})$ that specifies the linear weights at each point $\bm{x}$. Such weights describe a vector perpendicular to the decision boundary and reflect the local features’ importance at each specific location. Our model deals with survival data, which mostly comes as tabular data representing subjects’ features. Such features are usually explainable by themselves. This eliminates the role of the autoencoder part of the SENN model in extracting interpretable representation $\bm{h}(\bm{x})$, reducing it to the original feature space, i.e., $\bm{h}(\bm{x})=\bm{x}$, which is reflected in our model’s function shown in Eq. 13. An illustration of the CoxSE model is shown in Figure 3(a).
(a) CoxSE
(b) CoxSENAM
Figure 3: CoxSE and CoxSENAM models’ architecture
4.2 CoxSENAM: A Hybrid SENN + NAM Model
Based on the SENN and the NAM models, we also propose a hybrid model named CoxSENAM. As illustrated in Figure 3(b), the model has a structure similar to the CoxNAM model and a type of output and loss function similar to the CoxSE model.
The only difference between CoxSE and CoxSENAM is that the feature relevances are computed independently:
$f(\bm{x})=\sum_{x_{i}\in\bm{x}}w_{\theta_{i}}(x_{i})x_{i}$ (14)
where $w_{\theta_{i}}$ is the relevance function of the i’th feature ($x_{i}$) in the feature vector $\bm{x}$. The independence in computing feature relevances, although it constrains the model’s capacity to capture complex feature interactions, can contribute to mitigating the impact of noisy features—a benefit that will be demonstrated in subsequent analyses. On the other hand, this hybrid formulation enables the NAM model to leverage both regularization terms in the loss function, detailed in the following section, enhancing its stability and robustness.
4.3 Loss Function
The loss function for both CoxSE and CoxSENAM is composed of three terms. The first term is the Partial Likelihood function inherited from the CPH model (Eq. 15). This part of the loss function maximizes the concordance between the predicted hazards.
$\mathcal{L}_{1}(\theta)=\prod_{i=1}^{N}\frac{e^{f_{\theta}(\bm{x}_{i})}}{\sum_{\bm{x}_{j}\in R_{i}}e^{f_{\theta}(\bm{x}_{j})}}$ (15)
@@ 4.3 Loss Function
The second term (Eq. 16) minimizes the distance between $\nabla_{x}f_{\theta}(\bm{x})$, the gradient of $f_{\theta}(\bm{x})$ with respect to $\bm{x}$, and the predicted local weights $w_{\theta}(\bm{x})$:
$\mathcal{L}_{2}(\theta)=\|\nabla_{x}f_{\theta}(\bm{x})-w_{\theta}(\bm{x})\|_{2}$ (16)
This term is adopted from the SENN model to encourage explanation stability. A third term is added to regularize the learned weights and encourage the model to select a subset of the most important features while setting the weights of less important features to zero:
$\mathcal{L}_{3}(\theta)=\|w_{\theta}(\bm{x})\|_{1}$ (17)
The three terms comprise the complete loss function: